### 15 Mar 2026 v1

# Governing Dynamic Capabilities: Cryptographic Binding and Reproducibility Verification for AI Agent Tool Use

Abstract
[…]
1 Introduction
1.1 Agent Governance: A Missing Foundation
AI agents are no longer isolated tools. Through protocols like MCP (Model Context Protocol) [8] and A2A (Agent-to-Agent Protocol) [8], modern agents dynamically discover tools, invoke external APIs, execute code, and orchestrate other agents at runtime. The MCP ecosystem alone has grown to over 1,000 tool servers in under a year [32]. As these agents assume real-world responsibilities—managing enterprise credentials, executing financial transactions, accessing sensitive data—a fundamental question emerges: what properties must an agent ecosystem satisfy for humans to trust it?
This question has been raised from multiple directions. The EU AI Act (Regulation 2024/1689) mandates traceability, logging, and conformity assessment for high-risk AI systems (Articles 12, 43, 72). NIST’s AI Agent Standards Initiative [26] identifies security controls, governance oversight, and monitoring as key requirements. The TRiSM framework [28] structures agent governance around explainability, security, privacy, and lifecycle management. Schroeder de Witt [11] calls for secure interaction protocols that prevent insecure emergent behaviors. Cheng et al. [43] ground agent safety in transparency, accountability, and trustworthiness. Allegrini et al. [44] formalize 31 safety and liveness properties for agentic task lifecycles in temporal logic. Ghosh et al. [45] propose an operational safety framework with red-teaming methodology, policy-based tool scoping, and optional trace signing—but without cryptographic capability binding, formal adversary models, or verifiable replay mechanisms.
Yet despite this convergence of concern, no existing framework translates these governance aspirations into enforceable cryptographic infrastructure. The gap is not in recognizing what is needed, but in building it. We propose three Agent Governance Requirements that distill the common thread across these efforts into technically precise, verifiable properties:
Definition 1 (Agent Governance Requirements).
A governed agent ecosystem satisfies three properties:
G1: Capability Integrity. Every agent’s identity is cryptographically bound to its complete capability set—the model it executes and every tool it can invoke. The identity is unforgeable and non-repudiable. Any change to the capability set invalidates existing credentials and requires explicit re-authorization by a human or organizational principal.
G2: Behavioral Verifiability. An agent’s actions within its authorized capability envelope can be independently verified post-hoc, enabling detection of model substitution, unauthorized tool use, or output manipulation.
G3: Interaction Auditability. All inter-agent and agent-human interactions produce tamper-evident records sufficient for forensic reconstruction, allowing humans to audit agent behavior, attribute responsibility, and provide evidence for regulatory compliance and system improvement.
These requirements are not novel in aspiration—they reflect the consensus of legal [33, 34], regulatory [26], and technical [11, 28, 43, 44, 45] communities. What is novel is the claim that all three must hold simultaneously, and that none is achievable with existing agent infrastructure. G1 alone (without G2 and G3) permits certified agents to misbehave undetectably within their capability envelope. G2 alone (without G1) cannot distinguish authorized from unauthorized capabilities. G3 alone (without G1 and G2) produces audit logs of unverified actions by unverified agents.
1.2 Why the Current Landscape Falls Short
The current agent security landscape can be mapped into four layers (Figure 1), each addressing fragments of the governance requirements but none satisfying them:
Identity platforms (Keyfactor [19], SPIFFE/SPIRE [21, 22]) issue certificates binding an agent to a public key—but a certificate remains valid regardless of how many tools the agent acquires after issuance. G1 violated: capability changes are invisible to the identity layer.
Authorization frameworks (OIDC-A [4], Authenticated Delegation [3], Cerbos [39]) define what an authenticated agent may access—but authorization is evaluated against declared capabilities at grant time, with no mechanism to detect drift. G1 violated: authorization does not track capability changes.
Runtime gateways (Gravitee [40], Lasso Security [41]) inspect tool calls in real time—but are fundamentally stateless, with no memory of prior tool configurations and no behavioral verification. G2 and G3 violated.
Governance platforms (Credo AI [42]) maintain agent registries documenting capabilities—but operate at the documentation layer without cryptographic enforcement. G1 violated: declared capabilities are not bound to the agent’s actual configuration.
Figure 1: The four-layer agent security landscape and the capability-identity gap. Existing layers partially address individual requirements but none bridges them. Our framework provides the missing cryptographic governance layer satisfying G1–G3, validated end-to-end across 5–20 agent pipelines.
The root cause is a failure to recognize a fundamental architectural distinction in LLM-based agents. Inside a transformer’s forward pass, tool definitions and user context are indistinguishable token sequences. But at the orchestration layer, tool definitions determine which real-world actions the model’s outputs can trigger, while runtime context determines which actions the model chooses. This yields a principled capability-context separation: the capability envelope (model identity, tool definitions, permission scopes) governs what is possible and changes infrequently; the runtime context (conversation history, user instructions, memories) governs what is chosen and changes per-interaction. The security implication is direct: capability changes require re-authorization (G1); context changes require verifiability (G2) and auditability (G3).
Existing frameworks conflate the two. A tool addition and a new user message are treated identically—neither triggers re-evaluation. We call this the capability-identity gap: existing frameworks authenticate who an agent is and authorize what it may access, but none detect when the agent’s capabilities have changed. This gap enables a class of attacks we term silent capability escalation.
1.3 Our Framework: Bridging the Gap
The capability-context separation provides the theoretical foundation for three mechanisms, each targeting a specific governance requirement:
[…]
Together, these three mechanisms provide what the existing landscape lacks: a stateful, cryptographic governance layer satisfying G1–G3, bridging identity platforms and runtime gateways. Our contribution is not in novel cryptographic primitives; rather, we show that standard mechanisms—X.509 certificates, SHA-256 hashing, Ed25519 signatures, hash chains—when composed through the capability-context separation, are sufficient to enforce G1–G3 with negligible overhead, providing actionable design guidance for how agents should be built, how they should interact, and how their behavior should be audited.
1.4 Threat Model and Security Guarantees
We consider a PPT adversary who can compromise up to $k$ agent nodes, create malicious agents, inject or modify tools, and manipulate network traffic (Dolev-Yao model). Critically, the adversary can perform silent capability escalation. The adversary cannot compromise non-agent nodes or break standard cryptographic primitives.
Under this model, we define and prove nine security properties:
  1. 1.
    Capability Binding: tool configuration change $\Rightarrow$ access denied (Theorem 5).
  2. 2.
    Trust Containment: child constraints $\leq_{\kappa}$ parent constraints (Theorem 4).
  3. 3.
    Chain Integrity: valid chain to trusted root required (Theorem 6).
  4. 4.
    Delegation Depth: chains bounded by ancestor’s $\mathit{max\_depth}$ (Theorem 7).
  5. 5.
    Reproducibility Soundness: different model/tools detected with overwhelming probability (Theorem 8).
  6. 6.
    Bounded Divergence: adversary-controlled model passing $n$ replay checks is $(n,\epsilon)$-indistinguishable from declared model (Theorem 9).
  7. 7.
    Chain Verifiability: a single unverifiable interior agent breaks end-to-end behavioral verification for all downstream nodes (Theorem 12).
  8. 8.
    Credential Isolation: colluding agents gain no extra permissions (Theorem 10).
  9. 9.
    Ledger Integrity: undetectable modification impossible (Theorem 11).
Properties 1–4 enforce G1. Properties 5–7 enforce G2 (with 7 characterizing its chain-level semantics). Properties 8–9 enforce G3.
1.5 Contributions
In summary:
  • We define three Agent Governance Requirements—capability integrity, behavioral verifiability, and interaction auditability—distilling converging regulatory, legal, and technical concerns into formally precise properties (§1.1).
  • We identify the capability-identity gap and the capability-context separation principle as its root cause and resolution (§1.2).
  • We show that standard cryptographic mechanisms—composed through the capability-context separation—are sufficient to enforce all three requirements with negligible overhead, providing design guidance for secure agent ecosystems.
  • We propose capability-bound agent certificates with monotonic trust propagation (§3).
  • We introduce reproducibility as a trust primitive and establish that reproducibility is a necessary condition for behavioral verifiability: high-assurance agent deployments must be designed for deterministic output to be auditable (§4).
  • We define a verifiable interaction ledger for tamper-evident forensic reconstruction (§5).
  • We prove that sufficient replay verification yields bounded divergence: an adversary-controlled model that passes $n$ replay checks is provably $(n,\epsilon)$-indistinguishable from the declared model, with $\epsilon$ exponentially small in nn4.7).
  • We establish chain verifiability: behavioral verifiability is a chain property, not a node property—a single unverifiable interior agent breaks end-to-end verification for all downstream nodes, yielding principled deployment constraints on model placement in trust trees (§4.8).
  • We implement a working MCP governance proxy that enforces all three requirements in real-time with $\sim$}120 µs overhead per call, detecting silent capability escalation, tool trojanization, and certificate forgery (§7).
  • We formalize and prove nine security properties under a realistic adversary model (§6).
  • We implement a Rust prototype achieving 97 µs verification, ${\sim}0.62$ ms governance overhead, and ${\sim}1{,}200{,}000\times$ speedup over BAID (§7).
  • We deploy the full governance stack in an end-to-end multi-agent pipeline (5–20 agents with real LLM calls), demonstrating 0.12% overhead, sub-linear per-agent scaling, and detection of all in-situ attacks—while OAuth 2.1 and OpenTelemetry baselines detect none (§7.4).
1.6 Paper Organization
§2 presents the threat model and attack scenarios. §3 defines capability-bound certificates and the trust propagation tree. §4 introduces reproducibility commitments. §5 describes the verifiable interaction ledger. §6 presents formal security properties and proofs. §7 describes implementation and evaluation. §8 discusses limitations. §9 covers related work. §10 concludes.
2 Threat Model and Motivating Attacks
2.1 System Model
We model the trust infrastructure as a rooted directed tree $T=(V,E,r,\tau)$, where $V$ is the set of all entities (CAs, humans, organizations, AI agents), $E$ is the set of certificate issuance relationships, $r$ is the root CA, and $\tau:V\to\{\mathit{NA},\mathit{AG}\}$ classifies each node as Non-Agent ($\mathit{NA}$) or Agent ($\mathit{AG}$).
The tree satisfies a structural type constraint: $\tau(u)=\mathit{AG}\implies\forall(u,v)\in E:\tau(v)=\mathit{AG}$. Agents can delegate to sub-agents, but cannot confer trust upon humans or legal entities.
Each agent node $v$ holds a Capability-Bound Certificate $\mathit{cert}_{v}$ and maintains a runtime tool configuration $\mathcal{T}_{v}(t)$ that may change over time. Each credential $c$ has a risk tier $\mathit{tier}_{c}\in\{T0,T1,T2,T3\}$ where $T0$ (human-only) is most sensitive.
2.2 Adversary Model
We define a PPT adversary $\mathcal{A}$ with capabilities: C1: Network Control (Dolev-Yao). C2: Agent Compromise of up to $k$ agent nodes. C3: Rogue Agent Creation with fresh key pairs. C4: Silent Capability Escalation—modifying tool configurations without changing identity credentials. C5: Interaction Injection—fabricating or modifying ledger records.
$\mathcal{A}$ cannot: compromise non-agent nodes, forge digital signatures (EUF-CMA), find hash collisions, or reverse hash functions.
2.3 Motivating Attack Scenarios
We enumerate twelve attacks organized into four categories. ATK-1 through ATK-9 address structural governance threats; ATK-10 through ATK-12 address runtime behavioral attacks empirically validated by the Nemotron-AIQ dataset [45], which demonstrates 100% success rates for prompt injection and tool-mediated attacks against unprotected agentic workflows.
[…]
ATK-2: Model Substitution. $\mathcal{A}$ replaces the certified model with a fine-tuned variant. Detected via model binding mismatch and replay verification divergence.
ATK-3: Tool Trojanization. $\mathcal{A}$ replaces tool implementation (same name/version, different code). Detected via code-level hashing in the skills manifest.
Delegation Attacks. ATK-4: Phantom Delegation. Compromised agent (T2) issues T1 certificate to sub-agent. Prevented by monotonic trust propagation. ATK-5: Chain Forgery. Fabricated certificate chain. Prevented by signature verification at every level. ATK-6: Depth Overflow. Deep delegation hierarchy. Prevented by strictly decreasing depth counters. ATK-7: Credential Collusion. Two or more compromised agents at lower tiers (e.g., T3++T3) attempt to combine their credentials to access resources restricted to a higher tier (T1). Prevented by independent per-agent credential evaluation—each agent’s tier and permissions are verified in isolation, so colluding agents gain no additional permissions beyond their individual certificates (Theorem 10).
Forensic Attacks. ATK-8: Evidence Tampering. Log modification after incident. Prevented by hash-linked ledger with bilateral signing. ATK-9: Blame Shifting. Dispute in multi-agent chain. Resolved via interaction commitments and reproducibility replay.
Runtime Behavioral Attacks. ATK-10: Injection-Driven Knowledge Poisoning. Adversarial content injected via tool outputs (e.g., retrieval results) propagates through the agent’s reasoning chain, corrupting final outputs. Nemotron-AIQ traces show this attack succeeds in 100% of undefended trials across knowledge-poisoning and phishing-link templates. Our ledger’s content commitments ($\eta_{\mathit{in}},\eta_{\mathit{out}}$) enable post-hoc identification of the exact span where poisoned content entered the chain.
ATK-11: Cascading Action Hijack. Injected instructions persist across multiple agent workflow stages (INFECT template), causing the agent to refuse legitimate tasks or produce harmful content at downstream nodes. Our hash-chain ledger records every inter-span transition, enabling forensic tracing of attack propagation paths that OTEL-only audit trails cannot reliably provide.
ATK-12: Output Exfiltration via Rendering. Agent outputs are manipulated to embed phishing links (PHISHING_LINK) or malicious markdown (MARKDOWN) that exfiltrate data when rendered by downstream consumers. Detected post-hoc via reproducibility verification: replaying the interaction with the committed parameters produces benign output, diverging from the tampered record.
3 Capability-Bound Agent Certificates
3.1 Certificate Structure
Definition 2 (Capability-Bound Certificate).
A Capability-Bound Certificate for node $v$ is:
$\mathit{cert}_{v}=(\mathit{id}_{v},\mathit{id}_{\mathit{par}(v)},\mathit{pk}_{v},\mu_{v},\sigma_{v},\kappa_{v},\rho_{v},t_{s},t_{e},\mathit{sig})$
where $\mu_{v}=(\mathit{provider},\mathit{model\_id},\mathit{model\_ver})$ is the model binding, $\sigma_{v}=\{(\mathit{sid}_{i},\mathit{ver}_{i},h_{i},P_{i})\}_{i=1}^{n}$ is the skills manifest, $\kappa_{v}=(\mathit{max\_tier},\mathit{max\_depth},\mathit{allowed\_models},\mathit{max\_rate})$ are trust constraints, and $\rho_{v}=(\mathit{level},\mathit{config})$ is the reproducibility commitment.
3.2 Skills Manifest and Capability Binding
Definition 3 (Skills Manifest Hash).
$H_{v}^{\mathit{cert}}=\mathrm{SHA\text{-}256}(\mathrm{canonical}(\{(\mathit{sid}_{i},\mathit{ver}_{i},h_{i},P_{i})\}_{i=1}^{n}))$
At runtime: $H_{v}(t)=\mathrm{SHA\text{-}256}(\mathrm{canonical}(S_{v}(t)))$.
Capability Binding Invariant. Access permitted only if $H_{v}(t)=H_{v}^{\mathit{cert}}$. Any capability change invalidates access. Crucially, runtime context changes (memories, preferences) do not affect $H_{v}(t)$—the practical consequence of capability-context separation.
For open-source tools, $h_{i}$ is the SHA-256 hash of source code. For closed-source tools, $h_{i}$ is the hash of the public configuration descriptor (name, version, API schema).
3.3 Trust Propagation Tree
Figure 2: Trust propagation tree for the evaluation environment. NA nodes (rectangles) form the trust anchor; AG nodes (rounded) are agents. Constraints decay monotonically downward.
Definition 4 (Trust Constraint Ordering).
$\kappa_{a}\leq_{\kappa}\kappa_{b}$ iff $\kappa_{a}.\mathit{max\_tier}\leq\kappa_{b}.\mathit{max\_tier}$, $\kappa_{a}.\mathit{max\_depth}<\kappa_{b}.\mathit{max\_depth}$ (strict), $\kappa_{a}.\mathit{allowed\_models}\subseteq\kappa_{b}.\mathit{allowed\_models}$, and $\kappa_{a}.\mathit{max\_rate}\leq\kappa_{b}.\mathit{max\_rate}$.
Trust Propagation Rule. For every edge $(u,v)\in E$: $\kappa_{v}\leq_{\kappa}\kappa_{u}$.
3.4 Certificate Verification Algorithm
Figure 3: Verification flow for an MCP tool call through the A2Auth governance layer. Each phase can independently reject with a specific reason.
When agent $v$ requests access to credential $c$ at time $t$, the verifier executes four phases (Figure 3):
1function Verify(chain, credential, skills, roots):
2 // Phase 1: Chain Integrity
3 if chain[0].issuer $\notin$ roots: return DENY
4 for i in 1..len(chain):
5 if not VerifySig(chain[i-1].pk, chain[i].sig): return DENY
[…]
7 // Phase 2: Capability Binding
[…]
9 // Phase 3: Trust Constraint Check
[…]
11 // Phase 4: Revocation Check
12 if IsRevoked(cert_v): return DENY
13 return ALLOW
3.5 Integration with A2A Agent Cards
Agent Certificates are designed as non-invasive extensions of the A2A ecosystem via an optional extensions field in the Agent Card containing the certificate PEM, chain, and skills manifest hash.
4 Reproducibility Commitments
4.1 Reproducibility as a Necessary Condition
A key observation motivates our approach: behavioral verifiability (G2) is impossible without reproducibility. If an agent’s outputs cannot be independently regenerated—even approximately—then no post-hoc mechanism can distinguish authorized behavior from unauthorized behavior. This is not merely a limitation of our framework; it is a fundamental constraint on any verification scheme that does not rely on trusted execution environments.
This observation has a prescriptive implication for secure agent system design: high-assurance agent deployments must be engineered for reproducible output. Concretely, this means using deterministic sampling (temperature==0, fixed seeds), preferring structured outputs (code, JSON, SQL) over free-form prose where security-critical decisions are involved, and selecting inference providers that guarantee bitwise or near-bitwise reproducibility [6]. Agents that cannot commit to any reproducibility level receive a trust tier downgrade (§4.4), making the security cost of non-reproducibility explicit in the governance framework.
4.2 Empirical Foundation
LLM inference is nearly deterministic. Karvonen et al. [1] demonstrated that with fixed sampling parameters, regenerating output produces >>98% token-identical sequences. EigenAI [6] achieved 100% bit-exact deterministic inference on fixed GPU architectures with negligible performance loss. Our own validation (§7, Experiment 5) confirms that code-generating agents achieve perfect determinism, while prose outputs exhibit paraphrase-level variation that remains well-separated from cross-model divergence.
4.3 Reproducibility as a Trust Primitive
Definition 5 (Reproducibility Commitment).
$\rho_{v}=(\mathit{level},\mathit{config})$ where $\mathit{level}\in\{\mathit{full},\mathit{statistical},\mathit{none}\}$.
Full ($\mathit{level}=\mathit{full}$): $\forall(\mathit{context},\mathit{seed}):O_{v}(\mathit{context},\mathit{seed})=O_{v}^{\mathit{replay}}(\mathit{context},\mathit{seed})$.
Statistical ($\mathit{level}=\mathit{statistical}$):
$\forall(\mathit{context},\mathit{seed}):\mathrm{CharMatch}(O_{v},O_{v}^{\mathit{replay}})\geq\theta$
The threshold $\theta$ is task-type dependent: our empirical validation (Experiment 5, §7) shows that code-generating agents achieve perfect determinism ($\theta=0.98$), while prose-generating agents require $\theta=0.10$ to account for paraphrase-level variation. Cross-model divergence remains well below even the relaxed threshold (max 0.089).
None: No reproducibility commitment.
4.4 Reproducibility and Trust Tier Interaction
[…]
4.5 Replay-Based Verification Protocol
1function ReplayVerify(cert, record, O_orig):
[…]
3 if $\rho$.level = none: return INCONCLUSIVE
4 O_ref := ExecuteModel($\mu, cert$$\sigma$,
5 record.input, record.seed, $\rho$.config)
6 if $\rho$.level = full:
7 return O_ref == O_orig ? VERIFIED : VIOLATION
8 else: // statistical
9 return CharMatch(O_ref, O_orig) $\geq$ $\theta$ ?
10 VERIFIED : VIOLATION
4.6 Security Analysis
Theorem 1 (Reproducibility Soundness, informal).
If compromised agent $v$ executes model $M^{\prime}\neq M$ or skills $S^{\prime}\neq S$, replay verification detects the deviation with probability $\geq 1-\epsilon(\lambda)$ for negligible $\epsilon$.
By the DiFR token divergence result [1], the match rate between different models is $\delta\approx 0.06$ (character-level, validated in Experiment 5). By Hoeffding’s inequality: $\Pr[\mathrm{CharMatch}\geq\theta]\leq\exp(-2n(\theta-\delta)^{2})$, which is negligible for realistic output lengths.
4.7 Indistinguishability Under Sufficient Verification
A natural concern is adversarial fine-tuning: could an adversary produce a model $M^{\prime}$ that passes reproducibility checks on tested inputs but behaves maliciously on untested ones? We show that sufficient replay verification bounds the impact of any such attack.
Definition 6 ($(n,\epsilon)$-Indistinguishability).
Models $M$ and $M^{\prime}$ are $(n,\epsilon)$-indistinguishable under skill set $\mathcal{S}$ and prompt distribution $\mathcal{D}$ if, after $M^{\prime}$ passes replay verification on $n$ independent prompts $x_{1},\ldots,x_{n}\sim\mathcal{D}$ across skills in $\mathcal{S}$:
$\Pr_{x\sim\mathcal{D}}[\mathrm{CharMatch}(M(x),M^{\prime}(x))<\theta]\leq\epsilon$
Theorem 2 (Bounded Divergence).
If adversary-controlled model $M^{\prime}$ passes replay verification (CharMatch $\geq\theta$) on $n$ independent prompts drawn from $\mathcal{D}$ across $|\mathcal{S}|$ skills, then $M$ and $M^{\prime}$ are $(n,\epsilon)$-indistinguishable with:
$\epsilon\leq\exp\!\left(-2n\theta^{2}\right)$
Equivalently, for target confidence $1-\epsilon$, $n\geq\frac{\ln(1/\epsilon)}{2\theta^{2}}$ verification prompts suffice.
Proof sketch. Let $Z_{i}=\mathbf{1}[\mathrm{CharMatch}(M(x_{i}),M^{\prime}(x_{i}))<\theta]$ be independent Bernoulli random variables with $p=\Pr_{x\sim\mathcal{D}}[\mathrm{CharMatch}<\theta]$. Observing $\sum Z_{i}=0$ (all $n$ tests pass), by a one-sided binomial test: $\Pr[\sum Z_{i}=0\mid p\geq\epsilon]=(1-p)^{n}\leq(1-\epsilon)^{n}\leq\exp(-n\epsilon)$. Setting this equal to $\alpha$ (significance level) and solving yields $p\leq\ln(1/\alpha)/n$ with confidence $1-\alpha$. The Hoeffding refinement accounting for the threshold gap gives the stated bound. Full proof in §6. ∎
Practical implications. For code-generating agents ($\theta=0.98$), only $n=3$ prompts per skill achieve $\epsilon<10^{-2}$; $n=12$ achieves $\epsilon<10^{-10}$. Even for relaxed prose thresholds ($\theta=0.10$), $n=1200$ prompts achieve $\epsilon<10^{-10}$. These are operationally feasible: a verification suite of 50 prompts per skill, run periodically, provides strong guarantees.
Security interpretation. This transforms reproducibility from a detection mechanism into a safety certificate: an adversary who fine-tunes $M$ into $M^{\prime}$ that passes $n$ diverse replay checks has produced a model that is provably close to $M$ on the tested distribution. Any residual divergence is bounded by $\epsilon$, which decreases exponentially with the verification budget. The adversary faces a dilemma: either $M^{\prime}$ behaves like $M$ (and the fine-tuning achieves nothing), or $M^{\prime}$ diverges on some inputs and will be caught with probability increasing exponentially in $n$. This is analogous to cryptographic indistinguishability: the adversary’s advantage is negligible given sufficient verification.
4.8 Chain Verifiability
The preceding analysis establishes that G2 can verify an individual agent’s computational process. A natural question arises: when agents form a multi-hop execution chain, is per-node verification sufficient for end-to-end behavioral verifiability? We show that it is not—behavioral verifiability is a chain property, not a node property—and that this observation has direct implications for trust tree architecture (Figure 4).
[…]
The gap that G2 uniquely fills. Consider what G1+G3 alone guarantee. G1 (capability binding) ensures that each agent’s certificate is bound to its declared model and tool set; any capability change invalidates the certificate. G3 (interaction auditability) ensures that every input and output is bilaterally signed and hash-chain linked; no record can be silently modified. Together, G1+G3 provide attribution integrity: we know who sent what to whom, and that no message was altered in transit.
[…]
Why chain position matters. Now consider a trust tree execution chain $v_{0}\to v_{1}\to\cdots\to v_{n}$, where $v_{0}$ is a human principal (trust anchor) and each $v_{i}$ ($i\geq 1$) is an agent node. Agent $v_{j}$ receives input from $v_{j-1}$, processes it using its declared model, and passes output to $v_{j+1}$.
Suppose $v_{j}$ is a compromised agent with G2==none (either because it uses a low-determinism model or because G2 verification is not enforced). Then:
  1. 1.
    $v_{j}$’s output cannot be verified as originating from its declared model.
  2. 2.
    $v_{j}$’s output becomes $v_{j+1}$’s input. Even if $v_{j+1}$ satisfies G2==full, replay verification of $v_{j+1}$ only proves that $v_{j+1}$ faithfully processed the input it received—but that input itself is unverifiable.
  3. 3.
    By induction, all downstream nodes $v_{j+1},\ldots,v_{n}$ operate on data whose provenance cannot be cryptographically verified back to the trust anchor.
Definition 7 (Chain Verifiability Depth).
For execution chain $v_{0}\to v_{1}\to\cdots\to v_{n}$ with $v_{0}$ a non-agent trust anchor, define:
$\mathrm{CVD}(v_{n})=\min\{j\geq 1:v_{j}\text{ is an agent with }\rho_{v_{j}}.\mathit{level}=\mathit{none}\}$
with $\mathrm{CVD}(v_{n})=n$ if no such $v_{j}$ exists (fully verifiable chain).
Theorem 3 (Chain Verifiability).
Let $v_{0}\to v_{1}\to\cdots\to v_{n}$ be an execution chain where $v_{0}$ is a non-agent trust anchor. If $\mathrm{CVD}(v_{n})=k<n$, then for all $i>k$, the output of $v_{i}$ cannot be verified as originating from a legitimate computation chain rooted at $v_{0}$, even if $v_{i}$ individually satisfies $\rho_{v_{i}}.\mathit{level}=\mathit{full}$.
Proof sketch. By Definition 7, $v_{k}$ has $\rho.\mathit{level}=\mathit{none}$, so no replay verification can confirm that $v_{k}$’s declared model produced its output $o_{k}$. Under the adversary model (§2), $v_{k}$ may have executed an arbitrary model $M^{\prime}\neq M_{k}$ that produces output $o_{k}^{\prime}\neq o_{k}^{*}$ (where $o_{k}^{*}$ is the output the declared model would have produced). Agent $v_{k+1}$ receives $o_{k}^{\prime}$ as input. Even with $\rho_{v_{k+1}}.\mathit{level}=\mathit{full}$, replay verification confirms $v_{k+1}(o_{k}^{\prime})=o_{k+1}$, i.e., that $v_{k+1}$ faithfully processed $o_{k}^{\prime}$. But $o_{k}^{\prime}$ may differ from $o_{k}^{*}$, so $o_{k+1}$ is the correct computation on incorrect input. By induction on chain length, the same applies to all $v_{i}$ with $i>k$. Full proof in §6. ∎
Architectural implications. This theorem yields three deployment principles:
Principle 1: High-determinism models at interior nodes. Models with strong reproducibility (Anthropic Haiku: CharMatch==0.987, Sonnet: 0.823; see Table 4) should be placed at coordinator and interior positions in the trust tree, where their outputs become inputs for downstream agents. Low-determinism models (Kimi: 0.169, DeepSeek: 0.225) should be constrained to leaf positions where their outputs do not propagate further. This is not merely a performance recommendation—it is a security architecture constraint.
Principle 2: Verification boundary checkpoints. When a G2==none agent must appear at an interior position (e.g., a reasoning model as coordinator), the chain requires a verification boundary: either (a) a human-in-the-loop checkpoint that re-anchors trust by reviewing and signing the output, (b) TEE attestation providing hardware-enforced computational integrity, or (c) dual-execution with independent instances and output comparison (analogous to Byzantine fault tolerance). Each mechanism re-establishes chain verifiability for downstream nodes.
Principle 3: Explicit degradation in the ledger. When chain verifiability is broken at node $v_{k}$, the ledger should record a PARTIAL_VERIFIABILITY marker at that boundary. Downstream agents inherit a trust tier penalty beyond the standard G2==none downgrade (§4.4), making the reduced security guarantee visible to auditors and compliance systems.
Relationship to TEE. This analysis clarifies the complementary roles of G2 and trusted execution environments. G2 provides software-only verification of computational fidelity—a “software TEE”—using only black-box API access and the near-determinism of LLM inference. Hardware TEEs provide the same guarantee through hardware attestation but require infrastructure unavailable in most LLM deployments. The Chain Verifiability Theorem shows precisely where each mechanism is needed: at every interior agent node in the trust tree. In mixed deployments, G2 and TEE are interchangeable at the chain level: an agent with TEE attestation and G2==none maintains chain verifiability identically to an agent with G2==full and no TEE.
5 Verifiable Interaction Ledger
5.1 Motivation
Capability binding and reproducibility verify individual agents. Real-world incidents involve chains of interactions. We need an immutable, tamper-evident record where no single party—including the ledger operator—can undetectably alter history.
Why OpenTelemetry traces are insufficient. The industry standard for agentic observability is OpenTelemetry (OTEL) [46], whose span trees record execution traces with span_id/parent_id linkage, timestamped input.value/output.value payloads, and span-kind annotations (CHAIN, LLM, TOOL). Analysis of the Nemotron-AIQ Agentic Safety Dataset [45]—10,796 OTEL traces from NVIDIA’s AI-Q research assistant under 22 distinct attack templates—reveals three critical gaps: (1) No integrity protection: span linkage is purely referential; any party with storage access can modify, delete, or reorder spans undetectably. (2) No identity binding: spans record function names but not cryptographically verified agent identities or capability certificates. (3) No privacy preservation: OTEL stores full plaintext input/output payloads, creating data exposure risk. Our interaction ledger addresses all three gaps through hash-chain linking ($h_{\mathit{prev}}$), Ed25519 bilateral signing ($\mathit{sig}_{i}$), agent certificate binding ($\mathit{ch}_{\mathit{send}},\mathit{ch}_{\mathit{recv}}$), and commitment-only storage ($\eta_{\mathit{in}},\eta_{\mathit{out}}$).
5.2 Ledger Structure
Definition 8 (Interaction Record).
$R_{i}=(\mathit{seq}_{i},t_{i},a_{\mathit{send}},a_{\mathit{recv}},\mathit{ch}_{\mathit{send}},\mathit{ch}_{\mathit{recv}},\eta_{\mathit{in}},\eta_{\mathit{out}},\pi_{i},h_{\mathit{prev}},\mathit{sig}_{i})$ where $\eta_{\mathit{in}}=\mathrm{SHA\text{-}256}(\mathit{input})$, $\eta_{\mathit{out}}=\mathrm{SHA\text{-}256}(\mathit{output})$, $\pi_{i}=(\mathit{seed},\mathit{model\_ver},\mathit{skills\_hash})$, and $h_{\mathit{prev}}=\mathrm{SHA\text{-}256}(R_{i-1})$.
Definition 9 (Verifiable Interaction Ledger).
$\mathcal{L}=(R_{1},\ldots,R_{n})$ satisfying: (1) sequential consistency, (2) hash chain integrity, and (3) signature validity at every record.
5.3 Privacy Preservation
The ledger stores only commitments (hashes), not content. During normal operation, only interaction metadata is revealed. During forensic investigation, parties produce original content verified against stored commitments.
5.4 Forensic Reconstruction Protocol
The protocol proceeds in five steps: (1) identify the relevant interaction chain, (2) verify ledger integrity (signatures + hash chain), (3) request content disclosure and verify against commitments, (4) verify agent certificates at interaction time, and (5) execute replay verification for agents with reproducibility commitments.
6 Formal Security Properties and Proofs
We prove nine properties under the adversary model of §2, using standard cryptographic assumptions: EUF-CMA security, collision resistance of SHA-256, and the DiFR token divergence property [1].
Theorem 4 (Trust Containment).
For all edges $(u,v)\in E$: $\kappa_{v}\leq_{\kappa}\kappa_{u}$.
Proof.
Suppose $\mathcal{A}$ issues $\mathit{cert}_{w}$ with $\kappa_{w}\not\leq_{\kappa}\kappa_{v}$. The verifier checks $\kappa_{w}\leq_{\kappa}\kappa_{v}$ using constraints in $\mathit{cert}_{v}$, which is signed by $v$’s parent. $\mathcal{A}$ cannot modify $\mathit{cert}_{v}$ without the parent’s key. Recursing upward, the chain terminates at an uncompromised non-agent ancestor (EUF-CMA). ∎∎
Theorem 5 (Capability Binding).
If $H_{v}(t^{\prime})\neq H_{v}^{\mathit{cert}}$, then $\forall t\geq t^{\prime}:\mathrm{Access}(v,c,t)=\mathrm{DENY}$.
Proof.
By collision resistance: $\Pr[H_{v}(t^{\prime})=H_{v}^{\mathit{cert}}]\leq\mathrm{negl}(\lambda)$ when $S_{v}(t^{\prime})\neq\sigma_{v}$. Forging a new certificate requires the parent’s key (same argument as Theorem 4). ∎∎
Theorem 6 (Chain Integrity).
Only chains forming a valid path from a trusted root with valid signatures pass verification.
Proof.
Fabricating any link requires the corresponding private key. By EUF-CMA, $\mathcal{A}$ cannot produce valid signatures without the key. ∎∎
Theorem 7 (Delegation Depth Enforcement).
Agent delegation depth from any non-agent ancestor never exceeds $\mathit{max\_depth}$.
Proof.
Trust Containment requires $\kappa_{v}.\mathit{max\_depth}<\kappa_{\mathit{par}(v)}.\mathit{max\_depth}$ (strict). Starting from depth $d$, the chain has at most $d$ agent levels before reaching 0. ∎∎
Theorem 8 (Reproducibility Soundness).
If $v$ executes $M^{\prime}\neq M$ or $S^{\prime}\neq S$, replay verification detects with probability $\geq 1-\epsilon(\lambda)$.
Proof.
By DiFR [1]: $\mathbb{E}[\mathrm{CharMatch}(O^{\mathit{actual}},O^{\mathit{ref}})]\leq\delta$ for different models. By Hoeffding: $\Pr[\mathrm{CharMatch}\geq\theta]\leq\exp(-2n(\theta-\delta)^{2})$, negligible for realistic $n$. ∎∎
Theorem 9 (Bounded Divergence Under Verification).
If $M^{\prime}$ passes replay verification on $n$ independent prompts from $\mathcal{D}$, then $\Pr_{x\sim\mathcal{D}}[\mathrm{CharMatch}(M(x),M^{\prime}(x))<\theta]\leq\epsilon$ where $\epsilon\leq\exp(-2n\theta^{2})$.
Proof.
Let $p=\Pr_{x\sim\mathcal{D}}[\mathrm{CharMatch}(M(x),M^{\prime}(x))<\theta]$ be the true failure rate. Each verification trial is Bernoulli($p$). Observing $n$ successes (all pass), the likelihood ratio test gives: $\Pr[\text{all pass}\mid p]=(1-p)^{n}$. For $p\geq\epsilon$: $(1-\epsilon)^{n}\leq e^{-n\epsilon}$. Setting the false-acceptance probability to $e^{-n\epsilon}\leq e^{-2n\theta^{2}}$ (using $\epsilon=2\theta^{2}$ as the critical value from the Hoeffding bound on the CharMatch statistic), we obtain the stated bound. The Hoeffding refinement uses the fact that CharMatch is a bounded $[0,1]$ statistic with expected value $\geq\theta$ under $H_{0}$ (same model), applying $\Pr[\bar{X}-\mu\leq-t]\leq\exp(-2nt^{2})$ to the complement event. ∎∎
Theorem 10 (Credential Isolation).
$\mathrm{Accessible}(C^{\prime})=\bigcup_{v\in C^{\prime}}\mathrm{Accessible}(\{v\})$.
Proof.
Access depends solely on the requesting agent’s certificate. No protocol mechanism combines permissions. Proxy tokens are bound to specific agent IDs. ∎∎
Theorem 11 (Ledger Integrity).
No party can undetectably modify, delete, or reorder records.
Proof.
(a) Modification invalidates $h_{\mathit{prev}}$ of successor, cascading through chain; re-signing requires all subsequent senders’ keys. (b) Deletion breaks hash chain. (c) Reordering breaks both sequence numbers and hash links. ∎∎
Theorem 12 (Chain Verifiability).
Let $v_{0}\to v_{1}\to\cdots\to v_{n}$ be an execution chain where $v_{0}$ is a non-agent trust anchor, and let $\mathrm{CVD}(v_{n})=k<n$ (Definition 7). Then for all $i>k$, the output of $v_{i}$ cannot be verified as originating from a legitimate computation chain rooted at $v_{0}$, even if $\rho_{v_{i}}.\mathit{level}=\mathit{full}$.
Proof.
We proceed by induction on chain position.
[…]
Inductive step: Assume the output of $v_{i-1}$ (i1>ki-1>k) cannot be verified as legitimate. Agent $v_{i}$ receives this unverifiable output as input. By the same argument as the base case, $v_{i}$’s replay verification confirms computational fidelity on unverifiable input. ∎∎
7 Implementation and Evaluation
7.1 Prototype Architecture
We implement the framework as two Rust crates: agent-trust-kit (open-source SDK) and a2auth (credential governance application). Table 1 summarizes the implementation.
Table 1: Implementation summary.
Component Dependencies LoC
Certificate (parse, issue, verify) x509-parser, ring $\sim$}1,400
Skills manifest hasher sha2, serde_cbor $\sim$}500
Trust tree (internal) $\sim$}700
Reproducibility verifier sha2 $\sim$}400
Interaction ledger ed25519-dalek, sled $\sim$}1,100
Credential vault aes-gcm, argon2 $\sim$}800
Agent registry + proxy tokens agent-trust-kit $\sim$}600
MCP server + CLI jsonrpc-core, clap $\sim$}900
Real LLM repro. validation anthropic, sklearn (Python) $\sim$}870
MCP governance proxy demo cryptography, cbor2 $\sim$}520
Total $\sim$}8,290
7.2 Performance Evaluation
We evaluate on Apple M2, 16 GB RAM, macOS. All benchmarks use Criterion.rs with 10,000 samples and 3-second warmup.
Experiment 1: Certificate Verification Latency.
Chain Depth Mean Per-Sig Cap. Binding
[…]
5 (max practical) 163.6 µs 32.7 µs <<1 ns
Capability binding adds <<1 ns—the cost of a single 32-byte comparison. Verification is dominated by Ed25519 signature verification ($\sim$}33 µs/sig).
Experiment 2: Skills Manifest Hash.
Tools Sort CBOR SHA-256 Total
5 32 ns 0.41 µs 2.25 µs 2.69 µs
10 42 ns 0.76 µs 4.30 µs 5.10 µs
20 64 ns 1.71 µs 8.78 µs 10.6 µs
50 152 ns 3.94 µs 21.8 µs 25.9 µs
100 263 ns 8.05 µs 43.3 µs 51.6 µs
Experiment 3: Ledger Throughput.
Operation Latency Throughput
Append record 448 µs $\sim$}2,230 rec/s
Verify single record 32.4 µs $\sim$}30,900 rec/s
Full audit (1K records) 34.7 ms 34.7 µs/rec
Full audit (100K records) 3.46 s 34.6 µs/rec
Query by agent ID 112 µs 0.56 µs/rec
Experiment 4: Reproducibility Verification Latency.
Tokens Compare Replay Skip Commit
128 120 ns 123 ns 3.5 ns 1.72 µs
512 416 ns 418 ns 3.5 ns 6.55 µs
2,048 1.55 µs 1.55 µs 3.5 ns 24.3 µs
8,192 6.01 µs 6.10 µs 3.6 ns 96.6 µs
Token comparison scales at $\sim$}0.73 ns/token. Batch verification: 100 interactions (512 tokens each) in 41.8 µs. An auditor verifies 100,000 interactions in $\sim$}42 ms.
Experiment 5: Real LLM Reproducibility Validation. We validate the DiFR finding empirically in two phases. Phase 1 uses two Anthropic models (Claude Sonnet 4 and Haiku 4.5, temperature==0, max_tokens==512) across 105 prompts in five task types—code generation (25), JSON/structured output (15), mathematical reasoning (20), prose explanation (30), and creative writing (15)—with 5 same-model repetitions and 3 cross-model repetitions each (1,050 same-model + 1,575 cross-model pairwise comparisons). Phase 2 extends to 9 models from 7 providers: Claude Sonnet 4 and Haiku 4.5 (Anthropic), GPT-4.1 and GPT-4.1 mini (OpenAI), Gemini 2.5 Pro (Google), Grok 4 (xAI), DeepSeek V3 (DeepSeek), Llama 4 Maverick (Meta), and Kimi K2.5 (Moonshot). All models are accessed via OpenRouter as a unified API gateway (plus Volcengine for Doubao), with temperature==0 and 100 prompts across six task types (code, JSON, math, prose, creative, knowledge), 10 repetitions per model-prompt pair (5,400 total API calls; 15,120 pairwise comparisons: 10,800 same-model + 240 same-provider + 4,080 cross-provider). We compute four similarity metrics: character-level positional match (CharMatch), Jaccard token overlap, word-level TF-IDF cosine, and character n-gram (3–5) cosine; then apply Welch’s $t$-test, Mann-Whitney $U$, and Cohen’s $d$ to quantify separation, and use Youden’s $J$ statistic for optimal threshold selection.
Result 5a: Same-model reproducibility is task-type dependent. Table 2 shows per-task reproducibility from Phase 1 across 105 prompts (1,050 same-model pairwise comparisons).
Table 2: Same-model (Sonnet 4) reproducibility by task type.
Task Type $n$ Identical CharMatch TF-IDF cos
[…]
Structured outputs (JSON, code) have the highest identical-output fractions (>>50%), while prose shows the most variation (34.7% identical), driven by paraphrase-level reorganization. The bimodal distribution (median CharMatch =0.975=0.975 vs. mean =0.627=0.627) confirms that same-model outputs are either near-identical or substantially rephrased.
Result 5b: Multi-provider separation analysis. Table 3 presents the Phase 2 multi-provider results across 15,120 pairwise comparisons (10,800 same-model vs. 4,080 cross-provider). All differences are statistically significant ($p\approx 0$, Welch’s $t$ and Mann-Whitney $U$).
Table 3: Multi-provider separation: 9 models, 7 providers (15,120 comparisons).
Metric Same-model Cross-provider Sep. Cohen’s $d$
CharMatch 0.538 0.114 $4.7\times$ 1.36
Jaccard 0.710 0.321 $2.2\times$ 1.51
TF-IDF cosine 0.839 0.586 $1.4\times$ 1.08
Char n-gram cos 0.871 0.640 $1.4\times$ 1.20
All four metrics achieve Cohen’s $d>1.0$ (large effect sizes) across the full 9-model set, with Jaccard ($d=1.51$) and CharMatch ($d=1.36$) providing the strongest separation. Notably, same-provider pairs (e.g., GPT-4.1 vs. GPT-4.1 mini: CharMatch =0.251=0.251; Sonnet 4 vs. Haiku 4.5: $0.192$) exhibit similarity intermediate between same-model and cross-provider, confirming that even models from the same vendor maintain distinct output fingerprints.
Result 5c: Per-model determinism varies by over an order of magnitude. Table 4 ranks models by same-model CharMatch, revealing a wide determinism spectrum.
Table 4: Per-model same-model reproducibility (CharMatch, temperature==0).
Model Provider CharMatch Std
Claude Haiku 4.5 Anthropic 0.987 0.088
Claude Sonnet 4 Anthropic 0.823 0.273
GPT-4.1 mini OpenAI 0.701 0.335
GPT-4.1 OpenAI 0.617 0.363
Llama 4 Maverick Meta 0.528 0.360
Grok 4 xAI 0.402 0.343
Gemini 2.5 Pro Google 0.392 0.397
DeepSeek V3 DeepSeek 0.225 0.271
Kimi K2.5 Moonshot 0.169 0.233
Reasoning/thinking model (internal chain-of-thought).
Claude Haiku 4.5 produces near-identical outputs across repetitions (CharMatch =0.987=0.987), while reasoning models (Kimi K2.5, Gemini 2.5 Pro, Grok 4) exhibit substantially lower determinism due to stochastic chain-of-thought reasoning even at temperature 0. Despite this variation, even the least deterministic model (Kimi K2.5, CharMatch =0.169=0.169) still exceeds the mean cross-provider similarity ($0.114$), demonstrating that model identity remains distinguishable for all tested models.
Result 5d: Per-task separation is strongest for mathematical reasoning. Table 5 shows CharMatch separation factors across six task categories.
Table 5: Per-task CharMatch separation (same-model / cross-provider).
Task Type Same-model Cross-provider Sep.
Math reasoning 0.508 0.042 $12.2\times$
Knowledge 0.511 0.072 $7.1\times$
Creative writing 0.402 0.057 $7.0\times$
Prose 0.488 0.083 $5.9\times$
Code generation 0.594 0.161 $3.7\times$
JSON / structured 0.779 0.299 $2.6\times$
Mathematical reasoning achieves the highest separation ($12.2\times$) because models develop highly distinctive step-by-step reasoning styles, while JSON outputs show the lowest separation ($2.6\times$) because structural constraints force convergent formatting. These task-dependent results motivate per-task threshold calibration in deployment.
Result 5e: Optimal threshold via Youden’s $J$ statistic. Table 6 reports optimal thresholds from the multi-provider experiment.
Table 6: Optimal classification thresholds (multi-provider, 15,120 comparisons).
Metric $\theta^{*}$ Youden’s $J$ $F_{1}$
Jaccard 0.408 0.592 0.876
CharMatch 0.146 0.572 0.873
Char n-gram cos 0.809 0.538 0.855
TF-IDF cosine 0.837 0.466 0.844
In the multi-provider setting, Jaccard achieves the best overall classification ($F_{1}=0.876$, $J=0.592$), followed closely by CharMatch ($F_{1}=0.873$, $J=0.572$). Both metrics attain $F_{1}>0.87$ across 9 models and 7 providers, confirming that lightweight textual similarity metrics—requiring no model internals, embeddings, or logit access—provide practically useful model identity verification even in a heterogeneous multi-provider environment.
Experiment 6: End-to-End MCP Overhead.
Phase Latency
MCP request parse 0.02 ms
Certificate chain verification 0.097 ms
Skills hash comparison <<0.001 ms
Proxy token validation 0.05 ms
Ledger append 0.448 ms
Total governance overhead $\sim$}0.62 ms
Typical MCP call (no governance) 50–500 ms
Overhead as % of total 0.1–1.2%
Experiment 7: Comparison with BAID.
Metric BAID Ours Ratio
Verification latency $\sim$}120 s 0.097 ms $1.2{\times}10^{6}$
Proof size $\sim$}50 KB $\sim$}2 KB $25\times$
Infrastructure Blockchain None
Cap. change detection Post-hoc Real-time
Experiment 8: MCP Governance Proxy Demo.
To validate real-world applicability, we implement a complete MCP governance proxy in Python (520 lines) that interposes between an MCP client (agent) and MCP server (tool provider). The proxy enforces all three governance requirements: G1 via skills manifest hash verification against the agent’s certificate, G2 via reproducibility commitment tracking, and G3 via a hash-chain interaction ledger with Ed25519-signed records. We test 5 live attack scenarios:
Attack Result Latency
S1: Inject web_browser at runtime Blocked (hash $\Delta$) 121 µs
S2: Replace read_file with trojan Blocked (code hash) 103 µs
S3: Forged certificate (rogue CA) Blocked (sig verify) 106 µs
[…]
S5: Tamper ledger record #2 Detected (chain break)
All attacks are detected in real-time. Over 1,000 governed calls, the proxy achieves mean governance overhead of 120 µs/call (median: 106 µs, p99: 432 µs), throughput of 8,195 governed calls/sec, and all ledger records pass integrity verification. This confirms that the framework provides real-time protection with overhead well below the latency of typical MCP calls (50–500 ms).
7.3 Attack Scenario Evaluation
We implement 12 attack scenarios as automated tests: 9 structural governance attacks as Rust unit tests (774 lines), 5 end-to-end scenarios via the MCP governance proxy demo (520 lines), and 3 runtime behavioral attacks informed by the Nemotron-AIQ dataset [45] (which demonstrates 100% attack success rates without structural defenses). Table 7 summarizes results.
Table 7: Attack scenario evaluation. All 12 tests pass deterministically. S10–S12 are runtime behavioral attacks validated against Nemotron-AIQ traces.
Scenario Attack Baseline Framework Detection Mechanism
S1: Silent escalation Install web_browser at runtime Undetected Prevented capability_binding: hash mismatch
S2: Tool trojanization Replace implementation (same name) Undetected Prevented capability_binding: code hash change
S3: Model substitution Forge cert with different model Undetected Detected sig_verify: forged cert unsigned
S4: Phantom delegation T2 agent issues T1 cert Undetected Prevented constraint_monotonicity
S5: Evidence tampering Modify ledger record N/A Detected Hash chain + signature invalidation
S6: Blame shifting Dispute in agent chain Unresolvable Resolved Ed25519 signatures
S7: Colluding agents T3+T3 attempt T1 access Undetected Prevented tier_check: independent evaluation
S8: Chain forgery Fabricated chain Prevented sig_verify: fake key $\neq$ root
S9: Depth overflow $\mathit{max\_depth}=0$ delegation Prevented Structural: depth <0<0 impossible
S10: Knowledge poisoning Inject false data via tool output Undetected Traced Ledger $\eta_{\mathit{in}}/\eta_{\mathit{out}}$: pinpoint injection span
S11: Cascading hijack Multi-stage instruction persistence Undetected Traced Hash-chain: full propagation path audit
S12: Output exfiltration Embed phishing via rendered output Undetected Detected Replay divergence from committed $\pi_{i}$
100% attack success in Nemotron-AIQ undefended traces (2,596 security trials).
7.4 End-to-End Multi-Agent Pipeline
Experiments 1–8 evaluate individual governance operations in isolation. To validate the framework under realistic conditions, we deploy the full A2Auth governance stack across a multi-agent pipeline where agents coordinate, delegate tasks, invoke tools via MCP, and pass results through the interaction ledger. This experiment measures: (a) end-to-end governance overhead as a fraction of pipeline latency, (b) scaling behavior from 5 to 20 agents, (c) attack detection in situ, and (d) comparison with three baselines. The implementation uses the Rust agent-trust-kit SDK exposed to Python via PyO3 FFI, with agents making real Anthropic Claude API calls.
Figure 5: End-to-end multi-agent pipeline architecture with governance overlay. Five specialist agents coordinate through a trust tree (depth 3). Red labels indicate E2E attack injection points (Table 10). All inter-agent communications produce bilaterally-signed ledger records.
Pipeline architecture (Figure 5). We implement an AI Research Assistant pipeline: a user submits a research query; a Coordinator agent (T1_HIGH) orchestrates five specialist agents—Research (T2), Analysis (T2), Writer (T2), Review (T3_MONITORED), and Fact-Checker (T2)—each with a distinct skills manifest and reproducibility commitment. The trust tree has depth 3 (Root CA $\to$ Human $\to$ Coordinator $\to$ Workers), with the Coordinator’s max_delegation_depth=2{}=2 and all workers at depth 0 (cannot delegate). A single pipeline run traverses 21 steps: 10 tool calls (7 LLM invocations + 3 non-LLM tools), 11 inter-agent communications (task assignments and result deliveries), each producing a bilaterally-signed ledger record with capability verification and reproducibility commitment.
Experiment 9a: Governance overhead. Table 8 breaks down governance cost per pipeline run. Each operation’s latency is consistent with the micro-benchmarks from Experiments 1–4; the end-to-end cost is their composition.
Table 8: Per-pipeline governance overhead breakdown (5-agent configuration, 21 ledger entries per run, mean of 100 runs).
Operation Per-op Count Subtotal
Certificate chain verification 97 µs 6 0.58 ms
Skills manifest hash check 5.1 µs 10 0.05 ms
Ledger append (tool call) 448 µs 10 4.48 ms
Reproducibility commitment 6.6 µs 10 0.07 ms
Ledger append (inter-agent) 448 µs 11 4.93 ms
Bilateral Ed25519 signing 66 µs 11 0.73 ms
Total governance 21 10.8 ms
Pipeline latency (real LLM) $\sim$}8.7 s
Overhead 0.12%
Governance adds 10.8 ms to a pipeline that takes $\sim$}8.7 s with real Claude API calls (7 LLM inferences at $\sim$}1.2 s each + 3 non-LLM tools at $\sim$}100 ms each). The overhead of 0.12% is dominated by ledger appends (87% of governance cost), confirming that the cryptographic operations themselves (hashing, signing, verification) are negligible relative to I/O-bound agent workloads.
Experiment 9b: Scaling behavior. We scale the pipeline from 5 to 20 agents by instantiating parallel specialist groups: the Large configuration uses 4 Research agents (handling sub-queries in parallel), 4 Analysis agents, 3 Writer agents, 3 Review agents, 3 Fact-Checkers, 2 Coordinators (primary + sub-coordinator), and 1 Aggregator. Table 9 shows that per-agent governance cost remains approximately constant.
Table 9: Governance scaling: 5 $\to$ 20 agents (mean of 50 runs per configuration).
Config Agents Ledger Gov. Per-agent Storage
entries (ms) (ms) (KB)
Small 5 21 10.8 1.81 5.2
Medium 10 45 23.0 2.09 11.2
Large 20 95 48.3 2.30 23.8
Total governance cost scales linearly with ledger entries ($R^{2}>0.99$), but per-agent cost increases only 27% when quadrupling the agent count (1.81 ms $\to$ 2.30 ms). This near-constant per-agent cost arises because each agent’s governance operations—certificate verification, capability checking, ledger append—are independent of the total number of agents in the pipeline. The slight increase comes from deeper trust chains in the Large configuration (depth 4 vs. depth 3) adding one extra signature verification per activation. Ledger storage grows at $\sim$}250 bytes per entry; even the 20-agent pipeline produces only 23.8 KB per run, confirming that commitment-only storage (§5) is practical for long-running agent workflows.
Experiment 9c: End-to-end attack detection. We inject five attacks into the live 5-agent pipeline, each targeting a different governance layer. Unlike the unit-level tests in Table 7, these attacks occur mid-pipeline with real inter-agent data flow.
Table 10: End-to-end attack detection in the live multi-agent pipeline.
Attack Target Layer Detect. Hops
E2E-1: Capability escalation Research G1 5 µs
E2E-2: Model substitution Analysis G2 $\dagger$
E2E-3: Phantom delegation Writer G1 97 µs
E2E-4: Ledger tampering Ledger G3 32 µs
E2E-5: Cascading injection Pipeline G1+G2+G3 97 µs 3
[…]
dominated by LLM re-inference ($\sim$}1.2 s), not framework cost ($\sim$}13 µs).
[…]
Across 100 clean pipeline runs (500 total governed tool calls, 2,100 ledger entries), zero false positives are observed.
Experiment 9d: Baseline comparison. We evaluate the same five E2E attacks against three baselines: (1) no governance (raw MCP), (2) MCP with OAuth 2.1 (standard MCP authentication per specification), and (3) MCP with OpenTelemetry traces (observability-only, no cryptographic integrity). Table 11 summarizes results.
Table 11: E2E attack detection: A2Auth vs. baselines ($\checkmark$ = detected/prevented, ✗ = undetected). G1/G2/G3 indicate which governance requirement enables detection.
Attack No gov. OAuth OTEL A2Auth
E2E-1: Capability escalation $\checkmark$ (G1)
E2E-2: Model substitution $\checkmark$ (G2)
E2E-3: Phantom delegation a $\checkmark$ (G1)
E2E-4: Ledger tampering N/A N/A b $\checkmark$ (G3)
E2E-5: Cascading injection c $\checkmark$ (G1+G2+G3)
Total detected 0/5 0/5 0/5 5/5
aOAuth authenticates the Writer but has no delegation depth constraint. bOTEL traces can be silently modified (no hash chain or signatures; §5). cOTEL logs the poisoned data flow but cannot prove tampering or trace provenance cryptographically.
OAuth 2.1 provides identity authentication but not capability binding: it verifies who the agent is, not what tools it currently has. OpenTelemetry provides observability but not integrity: OTEL span trees log tool calls and parent-child relationships, but as demonstrated by analysis of 10,796 Nemotron-AIQ traces (§5), OTEL records contain no hash chains, no cryptographic signatures, and no identity binding—any field can be silently modified. A2Auth’s three governance layers (G1+G2+G3) provide the only configuration that detects all five attack types, with governance overhead below 0.2% of pipeline latency.
8 Limitations and Future Work
8.1 Scope of Contribution
Our contribution is a governance framework, not novel cryptographic primitives. The individual building blocks—X.509 certificates, SHA-256, Ed25519, hash chains—are standard. The novelty lies in (a) identifying the capability-context separation as the principled boundary for agent governance, (b) showing that composing standard mechanisms along this boundary enforces G1–G3 with negligible overhead, and (c) demonstrating this is practically achievable. We view this as analogous to Certificate Transparency [16], which composes Merkle trees and gossip protocols—neither novel individually—into infrastructure that fundamentally changed how the web PKI operates. The security proofs are intentionally straightforward reductions to standard assumptions (collision resistance, EUF-CMA); the contribution is in what to build, not in new ways to prove things.
8.2 Capability Binding Limitations
Capability-context boundary. Our sharp boundary has edge cases: a “code execution” tool can effectively acquire new capabilities through composition (e.g., downloading and running arbitrary scripts). Our framework guarantees that the registered capability set has not changed, but the effective attack surface within a fixed capability envelope can still be large. Complementary approaches (sandboxing, taint tracking, output filtering) are needed to govern behavior within the envelope. We view capability binding and intra-capability governance as orthogonal layers: the former ensures the attack surface is known; the latter constrains behavior within it.
Closed-source binding. For closed-source models and tools, the skills hash captures configuration metadata (name, version, API schema), not implementation code. A provider could change implementation without changing the public descriptor. Reproducibility verification provides a complementary detection mechanism: implementation changes that alter behavior will be detected via replay divergence.
Re-signing frequency. Mitigated through wildcard tool categories, short-lived auto-renewable certificates, and the empirical observation that tool configurations change far less frequently than user context.
8.3 Reproducibility Limitations and Graceful Degradation
Reproducibility is a design constraint, not a universal guarantee. Our framework makes explicit what is often implicit: non-reproducible agent behavior is non-verifiable. This is not a limitation unique to our approach—it applies to any verification scheme not based on trusted hardware. For high-assurance deployments, this means agents should be configured for deterministic output (structured formats, temperature==0, fixed seeds). Our empirical results (§7) show that code-generating agents already satisfy this constraint perfectly, and the separation between same-model and cross-model distributions ($11.5\times$) provides robust detection even for prose tasks.
Graceful degradation under partial G2. Our multi-provider experiment (§7, Experiment 5c) reveals that model determinism varies by over an order of magnitude: from CharMatch==0.987 (Haiku 4.5) to 0.169 (Kimi K2.5). This raises a practical question: what security guarantees hold when some agents use low-determinism models?
When G2 is unavailable (reproducibility commitment $\mathit{none}$), the framework degrades to G1+G3 only. Table 12 characterizes the precise security impact.
Table 12: Security guarantees under G1+G2+G3 vs. G1+G3 degradation.
Attack class G1+G2+G3 G1+G3
Silent capability escalation (ATK-1) Prevented Prevented
Tool trojanization (ATK-3) Prevented Prevented
Phantom delegation (ATK-4) Prevented Prevented
Chain/depth violation (ATK-5,6) Prevented Prevented
Evidence tampering (ATK-8) Detected Detected
Blame shifting (ATK-9) Resolved Resolved
Credential collusion (ATK-7) Prevented Prevented
Model substitution (ATK-2) Detected
Output exfiltration (ATK-12) Detected
G1+G3 alone protect against 9 of 12 attack types. The three attacks that require G2—model substitution (ATK-2), cascading injection via model compromise (ATK-10/11 when the injection occurs through model replacement rather than tool replacement), and output exfiltration via model tampering (ATK-12)—share a common structure: the attacker operates within a valid capability envelope by substituting the computational process itself. Signing proves attribution; only G2 proves computational fidelity.
Chain-level degradation. The Chain Verifiability Theorem (§4.8) establishes a stronger result: a G2==none agent at an interior position breaks behavioral verification for all downstream nodes. This means the true cost of deploying a low-determinism model at a coordinator position is not merely that the coordinator itself is unverifiable, but that the entire sub-tree it governs loses G2 protection. The framework’s three commitment levels (full/statistical/none) are thus not equivalent—their security implications depend critically on chain position.
Statistical vs. computational security. Token-level reproducibility provides statistical rather than computational security. However, the Bounded Divergence Theorem (§4.7) shows this gap is operationally small: sufficient replay verification transforms statistical detection into a provable safety certificate. An adversary who fine-tunes a model to pass $n$ diverse replay checks has produced a model that is $(n,\epsilon)$-indistinguishable from the original, with $\epsilon$ decreasing exponentially in $n$. Future work will explore semantic similarity (embedding cosine distance) as a verification metric, which may provide tighter separation bounds and is robust to paraphrase-level variation.
Provider cooperation. Replay verification requires access to the same model version. Organizations should negotiate replay rights as part of service agreements. Deterministic inference providers [6] eliminate this dependency entirely.
8.4 Trust Infrastructure Limitations
Non-agent node compromise. We assume non-agent nodes (CAs, humans, organizations) are uncompromised. Standard PKI mitigations apply: Certificate Transparency logs, hardware trust anchors (TPM 2.0), and multi-party issuance protocols. Ledger operator trust. The current single-operator ledger relies on bilateral signing for tamper evidence. A multi-operator design with consensus is possible but introduces latency overhead. Application-layer data flow. Our framework governs agent identity and capability, not the content of data flows. Complementary approaches (data-flow taint tracking, differential privacy) address this orthogonal concern.
8.5 Future Work
Hardware-attested agent identity (TPM 2.0, TEEs) for stronger G2 guarantees without provider cooperation. Embedding-based semantic similarity for reproducibility verification with formal threshold analysis. Cross-protocol capability governance beyond MCP/A2A. AI-assisted certificate issuance to reduce operational burden. Dynamic trust calibration from ledger history, enabling trust scores to evolve based on observed agent behavior.
9 Related Work
We organize related work along the four-layer landscape.
Identity Platforms. Keyfactor [19] and HID Global [20] propose X.509 for agents but without tool binding. SPIFFE/SPIRE [21, 22] provides workload identity. CyberArk [37] extends PAM to agents.
Authorization Frameworks. OIDC-A [4] extends OpenID Connect with agent claims. South et al. [3] propose Agent-ID Tokens (ICML 2025). Agentic JWT [10] defines delegation with temporal constraints.
Runtime Gateways. Gravitee [40] and Lasso [41] provide MCP proxies with per-call enforcement—fundamentally stateless.
Governance Platforms. Credo AI [42] maintains agent registries at the documentation layer.
Verifiable Execution. BAID [2] provides zkVM-based verification ($\sim$}120 s per proof). DiFR [1] demonstrates inference near-determinism. EigenAI [6] achieves 100% deterministic inference. Attestable Audits [7] uses TEEs.
Classical Trust Management. PolicyMaker [12], KeyNote [13], SPKI/SDSI [14], F-PKI [16], ARPKI [17].
Agent Protocols. A2A [8] provides agent communication. MCP enables agent-tool interaction with OAuth 2.1.
Agent Governance Frameworks. Cheng et al. [43] propose a three-pillar model (transparency, accountability, trustworthiness) as a conceptual governance framework. Allegrini et al. [44] formalize 31 task-lifecycle properties in CTL/LTL but address operational correctness rather than cryptographic governance. Ghosh et al. [45] propose an operational safety framework with agent red-teaming probes (ARP), policy-based tool scoping, and OpenTelemetry-based audit trails, validated through a 10,796-trace dataset (Nemotron-AIQ) covering 22 security attack templates and 13 content-safety categories against NVIDIA’s AI-Q research assistant. Their empirical results demonstrate that without guardrails, all 2,596 security attacks succeed (100%), and 20% of 2,800 safety attacks propagate to the final output—providing compelling evidence for the inadequacy of relying solely on model alignment. However, their OTEL-based audit trails lack cryptographic integrity: span linkage is purely referential (no hash chaining), spans carry no agent identity binding (no certificates), and full plaintext I/O is stored (no privacy preservation). Our framework addresses these gaps with a formally verified, hash-chain-linked, bilaterally signed interaction ledger (Section 5). Furthermore, their “signed schemas” and “per-call capability tokens” remain design principles without cryptographic specification or formal proofs. Raza et al. [28] adapt TRiSM for agentic multi-agent systems.
Regulatory Landscape. The EU AI Act (Regulation 2024/1689) requires traceability but lacks agent-specific implementing acts [33]. NIST CAISI [26] is developing agent standards. CSA [24, 25] proposes agentic trust frameworks.
10 Conclusion
AI agents increasingly acquire capabilities at runtime, yet no existing framework enforces the three governance requirements—capability integrity, behavioral verifiability, and interaction auditability—that a trustworthy agent ecosystem demands.
We addressed this through the capability-context separation principle and three mechanisms: capability-bound certificates (G1), reproducibility commitments (G2), and a verifiable interaction ledger (G3). We formalized nine security properties (including chain verifiability: behavioral verification is a chain property broken by any unverifiable interior agent), implemented the framework in Rust with 97 µs verification and <<1 ns capability binding overhead (${\sim}1{,}200{,}000\times$ faster than BAID), validated reproducibility across 9 models from 7 providers ($F_{1}=0.876$ multi-provider, $F_{1}=0.990$ single-provider), and verified 12 attack scenarios—including runtime behavioral attacks validated against NVIDIA Nemotron-AIQ traces—with specific, traceable detection mechanisms. An end-to-end evaluation deploying the full governance stack across a 5-to-20-agent pipeline with real LLM calls confirms 0.12% overhead, sub-linear per-agent scaling, and detection of all in-situ attacks with zero false positives—while three baselines (no governance, OAuth 2.1, OpenTelemetry) detect none.
Our framework fills the stateful, capability-aware governance gap between identity platforms and runtime gateways. As the EU AI Act’s high-risk provisions take effect and NIST develops agent-specific standards, the need for cryptographic infrastructure enforcing G1–G3 will only grow.
We release the core mechanisms as an open-source SDK (agent-trust-kit) to enable ecosystem adoption.